how do you reduce a fraction? i can't do the steps.
(63/6)(48/7)

Answer by stanbon(75887)

(63/6)(48/7)
You have to mentally or on paper factor the numbers.
= (7*3^2/2*3)(3*4^2/7)
Then you have to find factors that are common to numerator and denominator
like 3 and 7; cancel these because (3/3)= 1 and 7/7 = 1
-------------------
= (3/2)(3*2^4/1)
If you missed some common factor on the 1st go round do it again.
Cancel the common "2" that is in the numerator and the denominator:
= (3/1)(3*2^3/1)
When there is no more canceling to do, simplify the result:
= 9*8 = 72
=================
Cheers,
Stan H.